# Stratified Expert Cloning with Adaptive Selection for User Retention in Large-Scale Recommender Systems

### v1

@@ Abstract.
User retention has emerged as a critical challenge in large-scale recommender systems, significantly impacting the long-term success of online platforms. Existing methods often focus on short-term engagement metrics, failing to capture the complex dynamics of user preferences and behaviors over extended periods. While reinforcement learning (RL) approaches have shown promise in optimizing long-term rewards, they face difficulties in credit assignment, sample efficiency, and exploration when applied to the user retention problem. In this work, we propose Stratified Expert Cloning (SEC), a novel imitation learning framework that effectively leverages abundant logged data from high-retention users to learn robust recommendation policies. SEC introduces three key innovations: 1) a multi-level expert stratification strategy that captures the nuances in expert user behaviors at different retention levels; 2) an adaptive expert selection mechanism that dynamically assigns users to the most suitable policy based on their current state and historical retention level; and 3) an action entropy regularization technique that promotes recommendation diversity and mitigates the risk of policy collapse. Through extensive offline experiments and online A/B tests on two major video platforms, Kuaishou and Kuaishou Lite, with hundreds of millions of daily active users, we demonstrate SEC’s significant improvements over state-of-the-art methods in user retention. The results demonstrate significant improvements in user retention, with cumulative lifts of 0.098% and 0.122% in active days on Kuaishou and Kuaishou Lite respectively, additionally bringing several hundred thousand daily active users to each platform.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Figure 1. Schematic illustration of our recommendation system based on user state sequences. The system tracks user state transitions $(s_{i},s_{i+1},...)$ and makes content recommendations, i.e., action sequences $(a_{i},a_{i+1},...)$, with the ultimate goal of high user retention reward.
In the era of information abundance, recommender systems have become the cornerstone of online platforms, shaping user experiences and content consumption (Koren et al., 2009; Shi et al., 2017; Hidasi et al., 2015; Quadrana et al., 2018). Among the myriad of challenges faced by modern recommender systems, user retention has emerged as a critical metric that directly impacts long-term growth and sustainability. User retention refers to the ability to keep users consistently engaged and actively interacting with the platform over extended periods. However, optimizing for long-term user retention poses significant difficulties due to the complex dynamics of user preferences and behaviors that evolve over time.
Traditional recommender system approaches, such as collaborative filtering and content-based methods (He and Chua, 2017; Huang et al., 2016; Wang et al., 2019b, a), primarily focus on improving short-term engagement metrics, such as click-through rates or immediate user satisfaction. These approaches often fail to capture the temporal aspects of user interactions and the long-term effects of recommendations on user retention. Consequently, there is a growing need for recommendation strategies that prioritize user retention.
Recent advancements in reinforcement learning (RL) have shown promise in addressing the sequential nature of user interactions and optimizing for long-term rewards (Park and Han, 2013; Cai et al., 2023a; Liu et al., 2024b, a). RL-based methods formulate the recommendation problem as a Markov Decision Process (MDP) to learn policies that maximize cumulative rewards, thereby improving user retention. However, applying RL to optimize long-term user retention in real-world recommender systems remains challenging due to several factors.
Firstly, the delayed impact of recommendations on user retention introduces difficulties in credit assignment and reward attribution. The long-term effects of a recommendation may not be immediately observable, as shown in Figure 1, making it challenging for RL agents to associate specific actions with their corresponding long-term rewards. Secondly, the vast action space of all possible recommendations, coupled with the need for extensive exploration, can lead to sample inefficiency and suboptimal user experiences. Exploring irrelevant or poor-quality recommendations may frustrate users and hinder retention. Lastly, the complex and dynamic nature of user preferences and behaviors pose challenges in defining suitable reward functions that accurately capture the factors influencing user retention.
To address these challenges, we turn to imitation learning, a paradigm that has achieved notable success in large language model (Ouyang et al., 2022; Sun, 2023), robotics (Florence et al., 2022; Jia and Manocha, 2024) and autonomous driving (Samak et al., 2020; Yildirim et al., 2024) by learning from expert demonstrations. The key insight lies in leveraging abundant interaction data from high-retention users (treated as experts) to circumvent RL’s exploration challenges. By leveraging this valuable resource, we can effectively learn recommendation policies that directly optimize for long-term user retention while circumventing the difficulties faced by RL approaches.
Specifically, we propose Stratified Expert Cloning (SEC), a novel imitation learning framework for optimizing long-term user retention in recommender systems. SEC stratifies expert users into multiple levels based on their retention scores and learns specialized behavior cloning policies for each level. This hierarchical approach enables SEC to capture the nuances in expert user behaviors and preferences across different retention levels.
To adaptively select the most suitable expert policy for each user, SEC introduces an adaptive expert selection mechanism that considers the user’s current state and historical retention level. This mechanism dynamically assigns users to the appropriate expert level, ensuring that the recommendations align with the user’s evolving preferences and retention propensity.
Furthermore, SEC incorporates an action entropy regularization term to encourage recommendation diversity and mitigate the risk of overfitting to a narrow set of expert actions. By promoting exploration within the expert policy space, SEC enhances the generalization ability of the learned policies and improves long-term user engagement.
Through extensive offline experiments and large-scale online A/B tests on two major video platforms with hundreds of millions of daily active users, we demonstrate the significant improvements brought by SEC in user retention, compared to state-of-the-art methods. Our SEC framework offers a principled and practically effective approach for optimizing long-term user retention, providing recommender systems with a powerful tool to enhance user experiences and platform sustainability.
The main contributions of this work are summarized as follows:
@@
  • We propose SEC, a novel imitation learning framework for optimizing long-term user retention in recommender systems by learning hierarchical behavior cloning policies from expert users.
  • We introduce an adaptive expert selection mechanism that dynamically assigns users to the most suitable expert level based on their current state and historical interactions.
  • We incorporate an action entropy regularization term to encourage diverse and generalizable recommendation policies.
  • We conduct extensive offline experiments and online A/B tests on two major video platforms, demonstrating the superiority of SEC over state-of-the-art methods in improving user retention.
@@ 2.1. User Retention in Recommendation Systems
User retention is a crucial objective in recommendation systems, as long-term engagement often generates greater value than short-term interactions. Traditional methods like collaborative filtering and matrix factorization focused primarily on immediate engagement metrics such as click-through rates but struggled to model the temporal dynamics essential for user retention (Koren et al., 2009; Shi et al., 2017). To address this, recent frameworks like DT4Rec (Zhao et al., 2023b) leverage the Decision Transformer to optimize long-term user retention by modeling sequences of interactions. This approach incorporates an auto-discretized reward prompt that preserves partial order relationships between rewards, enhancing retention modeling over time. Other frameworks, such as IURO (Ding et al., 2023), explicitly model interpretable factors influencing retention, using contrastive learning to identify key content (”aha items”) responsible for sustained user engagement. Similarly, Generative Flow Networks (Liu et al., 2024a) have been applied to model retention by treating it as a probabilistic flow over user sessions, addressing the challenges of sparse and delayed retention signals.
In addition to these frameworks, research on recommendation diversity has shown that increasing content variety helps combat recommendation fatigue and improves user retention. Methods like Maximal Marginal Relevance (MMR) and other diversity-driven approaches (Carbonell and Goldstein, 1998; Park and Han, 2013) have demonstrated that diverse recommendations keep user interest high over time. Studies on customer purchase behavior (Kwon et al., 2020) in industries like fashion have empirically validated that diverse product recommendations not only increase purchase rates but also positively influence long-term retention, highlighting the effectiveness of recommendation diversity in both short-term and long-term engagement.
@@ 2.2. Imitation Learning
Reinforcement learning (RL) is effective for optimizing long-term rewards but often faces high sample complexity and the risk of suboptimal exploration, especially in recommendation systems where poor actions can harm user experiences. Imitation Learning (IL)  (Ng et al., 2000; Zare et al., 2024) addresses these challenges by training agents to mimic expert behaviors, avoiding the costly trial-and-error process of RL.
Among IL methods, Behavioral Cloning (BC) is particularly notable for its simplicity and efficiency (Pomerleau, 1991; Ross et al., 2011). Treating the problem as a supervised learning task allows BC to learn policies directly from expert demonstrations. This reduces the need for extensive interactions with the environment and lowers exploration risks. Behavior cloning has been effectively used in supervised fine-tuning and preference learning stages for Large Language Models  (Rafailov et al., 2024; Ethayarajh et al., 2024; Ouyang et al., 2022). Consequently, BC is particularly well-suited for applications that interact with users, such as recommender systems.
Advanced IL methods, such as Inverse Reinforcement Learning (IRL) (Ng et al., 2000) and Adversarial Imitation Learning (AIL) (Ho and Ermon, 2016), further extend IL’s capabilities. IRL infers reward functions behind expert behaviors, enabling generalization to unseen scenarios. Reinforcement Learning from Human Feedback (Ouyang et al., 2022; Wulfmeier et al., 2024) is a specialized form of IRL, where human feedback is treated as the expert guidance. AIL, exemplified by Generative Adversarial Imitation Learning (GAIL), aligns agent and expert distributions through adversarial training, excelling in continuous control tasks. IL’s versatility spans domains like large language model (Sun, 2023; Wulfmeier et al., 2024), robotics, autonomous driving, and game AI, with models like DQfD (Hester et al., 2018) and ChauffeurNet (Bansal et al., 2018) leveraging demonstrations to improve learning efficiency and generalization.
In recommendation systems, IL, particularly BC, holds promise for optimizing user retention by mimicking high-retention user behaviors. To enhance BC, we propose Stratified Expert Cloning (SEC), which stratifies expert users based on retention patterns and learns specialized policies for each group. This approach captures diverse user behaviors, improves long-term retention, and retains BC’s strengths in sample efficiency and interpretability.
@@ 3.2. Challenges in Optimizing Long-term User Retention
Traditional methods, such as collaborative filtering and content-based approaches (He and Chua, 2017; Huang et al., 2016; Wang et al., 2019b, a), primarily focus on short-term engagement metrics and fail to capture the long-term effects of recommendations on user retention. Reinforcement learning techniques (Park and Han, 2013; Cai et al., 2023a; Liu et al., 2024b, a), while promising, face difficulties in credit assignment, sample efficiency, and exploration when applied to the user retention problem.
@@ 3.3. Imitation Learning for User Retention
The SEC framework builds upon the strengths of imitation learning while addressing key challenges specific to recommender systems. By learning from high-retention experts in a principled manner, SEC provides a practical and effective approach for optimizing long-term user retention. The following sections detail the technical innovations of SEC and demonstrate its significant performance gains over existing methods.
@@ 4.1. Overview of the SEC Framework
Figure 3. T-SNE visualization of user state embeddings from the KuaiRand dataset, categorized by expert levels based on user return time. The plot shows the embedding space of level-1 experts, level-2 experts, level-3 experts, and non-expert users. Notably, some non-expert users align more closely with lower-level experts rather than the highest-retention group, motivating our multi-level stratification approach.
@@ 4.2.1. Multi-Level Expert Stratification
The SEC framework introduces a multi-level expert stratification approach to capture the nuances in expert user behaviors at different retention levels. Our key insight is that even among expert users with high overall retention, there exist distinct subgroups exhibiting different behavioral patterns. As shown in Figure 3, some non-expert users do not necessarily fall within the representation space of the highest-retention experts, but rather align more closely with the second or third level of experts. This observation motivates our design of a hierarchical expert stratification scheme.
@@ 5. Experiments
We evaluate the performance of our method SEC through extensive offline experiments and online A/B tests on two major video platforms with hundreds of millions of daily active users. During our offline testing, continuous actions are utilized, whereas discrete actions are used for the online experiments.
We aim to answer the following research questions through our experiments:
@@
  • RQ2: What are the effects of each component in SEC based on offline analysis?
@@
  • RQ4: Can SEC be extended to optimize other long-term values in online experiments?
@@ 5.1. Offline Experiments
Table 1. Overall Performance on KuaiRand-Pure dataset for different models.
Model Return Time Click Rate Long View Rate Like Rate
@@ 5.1.1. Experimental Setup
For offline evaluations, we utilize the KuaiRand dataset, a sequential recommendation dataset that provides an unbiased setting by incorporating random video exposures. KuaiRand encompasses a rich set of user feedback signals, including six positive signals (’click’, ’view time’, ’like’, ’comment’, ’follow’, and ’forward’) and two negative signals (’hate’ and ’leave’). To maintain focus and clarity in our analysis, we concentrate on the six positive signals and the two most relevant negative signals, while excluding less frequent feedback types. The KuaiRand dataset comprises interactions from 27,285 users on 7,551 items, totaling 1,436,609 interactions with a density of 0.70%.
To emulate realistic long-term user behavior in the context of short video recommendations, we employ the KuaiSim retention simulator (Zhao et al., 2023a). KuaiSim incorporates two primary modules: (1) a leave module that predicts the probability of a user exiting the current session, thus concluding an episode; and (2) a return module that estimates the daily likelihood of a user revisiting the platform, represented as a multinomial distribution. By leveraging KuaiSim, we can effectively assess the performance of recommendation algorithms in fostering long-term user engagement and retention. The dataset and KuaiSim retention configuration used in our experiments follow the settings described in the GFN model (Liu et al., 2024a).
In our offline experiments, We consider users with a return time less than or equal to 3 days as experts, and users with return times of one day, two days, and three days are considered experts at different levels. We employ the Policy network framework from GFN, which uses a Transformer architecture, for our behavior cloning approach. The user embeddings are pre-trained from the user model and fixed in our model. During inference, we set the number of clusters for each expert level to 1000 and define $\delta_{k}$ as half of the mean distance between clusters for each expert group. Optimization is carried out using the Adam optimizer with a batch size of 128. We apply a default balance parameter $\lambda$ of 0.01 and set the learning rate to 0.0003.
@@ 5.1.2. Baselines and Metrics
Our model is benchmarked against a diverse array of baselines, including:
  • CEM (Rubinstein and Kroese, 2004): A model-free optimization algorithm widely used in reinforcement learning tasks.
  • DIN (Zhou et al., 2018): A deep learning-based recommendation model that captures users’ dynamic interests from their historical behaviors.
  • TD3 (Fujimoto et al., 2018): An off-policy reinforcement learning algorithm that extends DDPG and addresses overestimation bias in Q-learning.
  • SAC (Haarnoja et al., 2018): An off-policy maximum entropy deep reinforcement learning algorithm that encourages exploration and robustness.
  • RLUR (Cai et al., 2023b): A reinforcement learning-based framework specifically designed for optimizing long-term user retention in recommender systems.
  • GFN (Liu et al., 2024a): A recently proposed generative model for sequential recommendation that learns a generative policy to model the probability distribution of user behavior sequences.
To evaluate the performance of the recommendation algorithms, we employ the following metrics:
  • Return Time: The average number of days it takes for a user to return to the platform after their last interaction. A lower return time indicates better user retention.
  • Click Rate: The ratio of the number of clicks to the total number of recommendations. A higher click rate suggests that the recommendations are more relevant to users’ interests.
  • Long View Rate: The proportion of video views that exceed a certain duration threshold (e.g., 50% of the video length). A higher long view rate implies that users find the recommended videos engaging and tend to watch them for a longer time.
  • Like Rate: The ratio of the number of likes to the total number of recommendations. A higher like rate indicates that users express more immediate satisfaction with the recommended items.
These metrics provide a comprehensive assessment of the recommendation algorithms’ performance in terms of user retention (Return Time), user engagement (Click Rate and Long View Rate), and user satisfaction (Like Rate). By evaluating the models across these diverse metrics, we can gain insights into their effectiveness in optimizing long-term user retention and delivering a satisfactory user experience.
@@
Variant Return Day Click Rate Long View Rate

### v2

# Stratified Expert Cloning for Retention-Aware Recommendation at Scale

@@ Abstract.
User retention is critical in large-scale recommender systems, significantly influencing online platforms’ long-term success. Existing methods typically focus on short-term engagement, neglecting the evolving dynamics of user behaviors over time. Reinforcement learning (RL) methods, though promising for optimizing long-term rewards, face challenges like delayed credit assignment and sample inefficiency.
We introduce Stratified Expert Cloning (SEC), an imitation learning framework that leverages abundant interaction data from high-retention users to learn robust policies. SEC incorporates: 1) multi-level expert stratification to model diverse retention behaviors; 2) adaptive expert selection to dynamically match users with appropriate policies based on their state and retention history; and 3) action entropy regularization to enhance recommendation diversity and policy generalization.
Extensive offline evaluations and online A/B tests on major video platforms (Kuaishou and Kuaishou Lite) with hundreds of millions of users validate SEC’s effectiveness. Results show substantial improvements, achieving cumulative lifts of 0.098% and 0.122% in active days on the two platforms respectively, each translating into over 200,000 additional daily active users.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: xx/xxisbn: 979-8-4007-2040-6/2025/11ccs: Information systems Recommender systems
@@ 1. Introduction
[…]
In the era of information abundance, recommender systems are central to online platforms, significantly influencing user experience and content consumption (Koren et al., 2009; Shi et al., 2017; Hidasi et al., 2015; Quadrana et al., 2018). Among the many challenges faced by recommender systems, user retention—defined as sustaining long-term user engagement—is critical for sustainable growth. However, it remains difficult due to the constantly evolving nature of user preferences and behaviors, as illustrated in Figure 1.
Traditional methods such as collaborative filtering and content-based approaches (He and Chua, 2017; Huang et al., 2016; Wang et al., 2019b, a) focus on short-term engagement metrics like click-through rate, often overlooking the dynamics of long-term user retention. While recent advances in reinforcement learning (RL) (Park and Han, 2013; Cai et al., 2023a; Liu et al., 2024b, a) offer promising solutions for sequential decision-making in recommender systems, they still face significant challenges—including delayed rewards, large action spaces, and evolving user behaviors—that hinder effective credit assignment, sample efficiency, and reward design.
To address these limitations, we turn to imitation learning, leveraging expert demonstration data—particularly from high-retention users—to bypass the inefficiencies of RL exploration. We propose Stratified Expert Cloning (SEC), an imitation learning framework that stratifies expert users by retention levels and learns specialized behavior policies tailored to each group.
SEC encompasses three core technical innovations designed to optimize long-term user retention: (1) multi-level expert stratification, which segments expert users into distinct retention-based hierarchies and learns specialized recommendation policies for each level, enabling nuanced behavioral modeling across the retention spectrum; (2) adaptive expert selection, which dynamically assigns users to the most appropriate expert policy based on their current state embedding and historical interaction patterns, providing personalized recommendation strategies; and (3) action entropy regularization, which promotes recommendation diversity through nuclear norm maximization in the action space, preventing policy collapse while enhancing exploration and generalization across user interests. Together, these components form a comprehensive framework that effectively leverages expert behaviors to improve user retention in large-scale recommender systems.
Through comprehensive offline and extensive online A/B tests on two major video platforms, SEC substantially improves user retention compared to state-of-the-art methods, highlighting its effectiveness and practicality in large-scale recommendation scenarios.
The contributions of our work are summarized as follows:
@@
  • We propose SEC, a novel imitation learning framework for hierarchical behavior cloning from expert users to optimize long-term retention.
  • We introduce adaptive expert selection based on user state and history to personalize recommendations.
  • We employ action entropy regularization to enhance recommendation diversity and generalization.
  • We validate SEC’s superiority through extensive experiments and online evaluations on large-scale platforms.
@@ 2.1. User Retention in Recommendation Systems
User retention is crucial for recommender systems as long-term engagement generates higher value than short-term interactions. Traditional methods such as collaborative filtering and matrix factorization primarily focus on immediate engagement metrics (e.g., click-through rates) but often struggle with modeling temporal dynamics necessary for long-term retention (Koren et al., 2009; Shi et al., 2017). Recent advances like DT4Rec (Zhao et al., 2023b) utilize Decision Transformers to optimize retention by modeling interaction sequences with discretized reward prompts. Similarly, IURO (Ding et al., 2023) explicitly models interpretable factors influencing retention, while Generative Flow Networks (Liu et al., 2024a) treat user retention as probabilistic flows over sessions, effectively handling sparse and delayed signals.
Moreover, recommendation diversity methods like Maximal Marginal Relevance (MMR) (Carbonell and Goldstein, 1998; Park and Han, 2013) demonstrate that diversified recommendations help maintain user interest, reducing recommendation fatigue and improving retention. Empirical studies (Kwon et al., 2020) confirm diversity significantly boosts long-term engagement and purchase rates.
@@ 2.2. Imitation Learning
Reinforcement Learning (RL), though effective for long-term optimization, faces challenges like high sample complexity and risks from exploration, particularly in recommender systems. Imitation Learning (IL) (Ng et al., 2000; Zare et al., 2024) addresses these issues by learning from expert behaviors, eliminating costly trial-and-error interactions.
Behavioral Cloning (BC), an IL method notable for its simplicity, learns directly from expert demonstrations as supervised learning (Pomerleau, 1991; Ross et al., 2011). BC has proven effective in diverse domains, including large language model (Ouyang et al., 2022; Rafailov et al., 2024; Ethayarajh et al., 2024), robotics (Florence et al., 2022; Jia and Manocha, 2024) and autonomous driving (Samak et al., 2020; Yildirim et al., 2024).
Advanced IL methods like Inverse Reinforcement Learning (IRL) (Ng et al., 2000) and Generative Adversarial Imitation Learning (GAIL) (Ho and Ermon, 2016) infer reward functions and align agent-expert distributions respectively, enhancing IL’s generalizability. Models such as DQfD (Hester et al., 2018) and ChauffeurNet (Bansal et al., 2018) exemplify IL’s versatility across complex tasks.
In recommendation contexts, IL (particularly BC) is promising for optimizing user retention by replicating behaviors of high-retention users. Our proposed Stratified Expert Cloning (SEC) framework enhances BC by learning specialized policies from stratified expert users, effectively capturing diverse behaviors and promoting sustained user engagement.
@@ 3.2. Challenges in Optimizing Long-term User Retention
Traditional methods, such as collaborative filtering and content-based approaches (He and Chua, 2017; Huang et al., 2016; Wang et al., 2019b, a), primarily focus on short-term engagement metrics and fail to capture the long-term effects of recommendations on user retention. Reinforcement learning techniques (Park and Han, 2013; Cai et al., 2023a; Liu et al., 2024b, a), while promising, face difficulties in credit assignment, sample efficiency, reward definition and exploration when applied to the user retention problem.
@@ 3.3. Imitation Learning for User Retention
Based on these insights, we propose a novel imitation learning framework specifically designed to address the challenges of optimizing user retention in recommender systems. The following section details our approach.
@@ 4.1. Overview of the SEC Framework
Figure 3. T-SNE visualization of user state embeddings from the KuaiRand dataset, categorized by expert levels based on retention patterns. The visualization shows distinct clusters for level-1, level-2, and level-3 experts, with non-expert users distributed across the space. Notably, many non-expert users align more closely with lower-level experts (level-2 and level-3) rather than with the highest-retention experts (level-1), demonstrating that user behaviors exist on a continuum of retention patterns. This empirical finding strongly supports our multi-level expert stratification approach over traditional binary expert/non-expert classifications.
@@ 4.2.1. Multi-Level Expert Stratification
The SEC framework introduces a multi-level expert stratification approach to capture the nuances in expert user behaviors at different retention levels. Our key insight is that even among expert users with high overall retention, there exist distinct subgroups exhibiting different behavioral patterns. As illustrated in Figure 3, our analysis reveals that non-expert users distribute heterogeneously in the representation space, with many aligning more closely with level-2 or level-3 experts rather than with the highest-retention experts (level-1). This embedding pattern demonstrates that user retention behaviors exist on a continuum rather than in binary states, providing strong empirical evidence for our hierarchical expert stratification approach. By learning specialized policies from multiple expert levels, we can more effectively guide diverse users toward improved retention through personalized recommendation strategies.
@@ 5. Experiments
We evaluate SEC through offline experiments and online A/B tests on two major video platforms with hundreds of millions of daily active users. During our offline testing, continuous actions are utilized, whereas discrete actions are used for the online experiments. We aim to answer the following research questions:
@@
  • RQ2: What are the effects of each component in SEC?
@@
  • RQ4: Can SEC be extended to optimize other long-term values?
@@ 5.1. Offline Experiments
Table 1. Overall Performance on KuaiRand-Pure dataset.
Model Return Time $\downarrow$ Click Rate $\uparrow$ Long View Rate $\uparrow$ Like Rate $\uparrow$
@@ 5.1.1. Experimental Setup
For offline evaluations, we utilize the KuaiRand dataset, a sequential recommendation dataset with random video exposures. It includes user feedback signals such as ’click’, ’view time’, ’like’, ’comment’, ’follow’, ’forward’, ’hate’, and ’leave’. The dataset comprises interactions from 27,285 users on 7,551 items (1,436,609 interactions, 0.70% density).
We employ the KuaiSim retention simulator (Zhao et al., 2023a) to emulate long-term user behavior, which includes a leave module predicting session exits and a return module estimating platform revisit likelihood. The configurations follow the GFN model settings (Liu et al., 2024a).
In our experiments, users with return times of ¡=3 days are considered experts at different levels. We use the Policy network framework from GFN with a Transformer architecture. User embeddings are pre-trained and fixed. During inference, we set 1000 clusters per expert level with $\delta_{k}$ as half the mean inter-cluster distance. Optimization uses Adam with batch size 128, $\lambda=0.01$, and learning rate 0.0003.
@@ 5.1.2. Baselines and Metrics
We benchmark against CEM (Rubinstein and Kroese, 2004), DIN (Zhou et al., 2018), TD3 (Fujimoto et al., 2018), SAC (Haarnoja et al., 2018), RLUR (Cai et al., 2023b), and GFN (Liu et al., 2024a). We evaluate using:
  • Return Time: Average days until platform return (lower is better)
  • Click Rate: Clicks per recommendation (higher is better)
  • Long View Rate: Views exceeding duration threshold (higher is better)
  • Like Rate: Likes per recommendation (higher is better)
@@
[…]